OpenStudy (anonymous):

Need help with this problem pleasse... For each function, what is the output for the given input? Forf(x)=4x+7, find f(4).

OpenStudy (mayankdevnani):

answer is 23

OpenStudy (mayankdevnani):

@barrelracer23

OpenStudy (anonymous):

Just replace x with 4 in 4x + 7 4(4) + 7 = 16 + 7 = 23.

OpenStudy (anonymous):

@mayankdevnani please don't give the direct answer, give the explanation too.

OpenStudy (anonymous):

@pratu043 Thank you so much:)